# Softpick: No Attention Sink, No Massive Activations with Rectified Softmax

### v1

@@ Appendix C Full Result on Quantization
Table 9: Comparison of softpick vs softmax performance for HQQ Badri and Shaji [2023] quantization methods. $\uparrow$=Higher is Better, $\downarrow$=Lower is Better. $\Delta$ = softpick - softmax.
Task Metric Quantization Softmax Softpick $\Delta$ (S - V)
Arc Easy Acc Norm $\uparrow$ 2-bit 26.64 36.53 +9.89
Acc $\uparrow$ 2-bit 26.77 37.37 +10.60
Acc Norm $\uparrow$ 3-bit 48.61 52.44 +3.83
Acc $\uparrow$ 3-bit 52.78 58.12 +5.34
Lambada Acc $\uparrow$ 2-bit 0.00 5.71 +5.71
Perplexity $\downarrow$ 2-bit 21819832.38 3745.16 -21816087.22
Acc $\uparrow$ 3-bit 26.28 32.19 +5.91
Perplexity $\downarrow$ 3-bit 76.73 41.80 -34.93
Piqa Acc Norm $\uparrow$ 2-bit 50.54 55.28 +4.74
Acc $\uparrow$ 2-bit 52.39 56.15 +3.76
Acc Norm $\uparrow$ 3-bit 62.40 65.23 +2.83
Acc $\uparrow$ 3-bit 63.17 65.94 +2.77
Sciq Acc Norm $\uparrow$ 2-bit 22.00 51.20 +29.20
Acc $\uparrow$ 2-bit 21.40 53.50 +32.10
Acc Norm $\uparrow$ 3-bit 71.50 71.90 +0.40
Acc $\uparrow$ 3-bit 80.50 80.50 +0.00
Wikitext Word Perplexity $\downarrow$ 2-bit 151861.92 5853.90 -146008.03
3-bit 288.09 202.28 -85.81
Appendix D More Attention Maps
@@
Figure 7: More attention maps of the softmax model on sample text input 1.
Figure 8: More attention maps of the softpick model on sample text input 1.
Figure 9: More attention maps of the softmax model on sample text input 2.
Figure 10: More attention maps of the softpick model on sample text input 2.
Figure 11: More attention maps of the softmax model on sample text input 3.
Figure 12: More attention maps of the softpick model on sample text input 3.

### v2

@@ Appendix C Full Result on Quantization
Table 10: Comparison of softpick vs softmax performance for HQQ [2] quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Quantization Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 2-bit 26.68 36.36 +9.68
Acc $\uparrow$ 2-bit 26.81 37.54 +10.73
Acc Norm $\uparrow$ 3-bit 48.65 52.57 +3.91
Acc $\uparrow$ 3-bit 52.78 57.95 +5.18
Lambada Acc $\uparrow$ 2-bit 0.00 5.76 +5.76
Perplexity $\downarrow$ 2-bit 21857842.67 3741.81 -21854100.86
Acc $\uparrow$ 3-bit 26.18 32.14 +5.96
Perplexity $\downarrow$ 3-bit 76.81 41.85 -34.96
Piqa Acc Norm $\uparrow$ 2-bit 50.38 55.11 +4.73
Acc $\uparrow$ 2-bit 52.29 56.04 +3.75
Acc Norm $\uparrow$ 3-bit 62.40 65.78 +3.37
Acc $\uparrow$ 3-bit 63.55 66.00 +2.45
Sciq Acc Norm $\uparrow$ 2-bit 22.20 51.40 +29.20
Acc $\uparrow$ 2-bit 21.60 53.60 +32.00
Acc Norm $\uparrow$ 3-bit 71.60 72.10 +0.50
Acc $\uparrow$ 3-bit 80.50 80.70 +0.20
Wikitext Word Perplexity $\downarrow$ 2-bit 127870.20 2139.86 -125730.34
3-bit 35.52 30.24 -5.29
Appendix D Full Result on Quantization on 1.8B
Table 11: Comparison of softpick vs softmax performance for BNB [13, 14] quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Quantization Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 4-bit 66.75 58.84 -7.91
Acc $\uparrow$ 4-bit 71.93 67.05 -4.88
Acc Norm $\uparrow$ 8-bit 67.26 62.21 -5.05
Acc $\uparrow$ 8-bit 72.22 69.07 -3.16
Lambada Acc $\uparrow$ 4-bit 47.84 41.26 -6.58
Perplexity $\downarrow$ 4-bit 12.90 18.56 +5.66
Acc $\uparrow$ 8-bit 48.65 43.14 -5.51
Perplexity $\downarrow$ 8-bit 11.74 16.11 +4.36
Piqa Acc Norm $\uparrow$ 4-bit 72.14 70.57 -1.58
Acc $\uparrow$ 4-bit 72.63 71.00 -1.63
Acc Norm $\uparrow$ 8-bit 72.96 70.40 -2.56
Acc $\uparrow$ 8-bit 73.56 71.22 -2.34
Sciq Acc Norm $\uparrow$ 4-bit 86.90 79.40 -7.50
Acc $\uparrow$ 4-bit 90.90 86.90 -4.00
Acc Norm $\uparrow$ 8-bit 86.00 79.90 -6.10
Acc $\uparrow$ 8-bit 90.30 86.90 -3.40
Wikitext Word Perplexity $\downarrow$ 4-bit 16.17 18.97 +2.80
8-bit 15.20 17.94 +2.74
Table 12: Comparison of softpick vs softmax performance for GPTQ [17] quantization. $\uparrow$=Higher is Better, $\downarrow$=Lower is Better. $\Delta$ = Softpick - Softmax.
Task Metric Quantization Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 2-bit 28.62 34.60 +5.98
Acc $\uparrow$ 2-bit 29.67 33.88 +4.21
Acc Norm $\uparrow$ 3-bit 61.28 57.49 -3.79
Acc $\uparrow$ 3-bit 64.90 64.27 -0.63
Acc Norm $\uparrow$ 4-bit 65.74 61.87 -3.87
Acc $\uparrow$ 4-bit 70.20 68.73 -1.47
Acc Norm $\uparrow$ 8-bit 66.96 61.95 -5.01
Acc $\uparrow$ 8-bit 72.77 68.60 -4.17
Lambada Acc $\uparrow$ 2-bit 1.14 1.59 +0.45
Perplexity $\downarrow$ 2-bit 27256.52 18730.23 -8526.29
Acc $\uparrow$ 3-bit 40.71 36.77 -3.94
Perplexity $\downarrow$ 3-bit 20.58 25.33 +4.75
Acc $\uparrow$ 4-bit 49.64 40.95 -8.69
Perplexity $\downarrow$ 4-bit 11.43 17.94 +6.50
Acc $\uparrow$ 8-bit 49.60 43.41 -6.19
Perplexity $\downarrow$ 8-bit 11.35 15.82 +4.47
Piqa Acc Norm $\uparrow$ 2-bit 53.97 53.92 -0.05
Acc $\uparrow$ 2-bit 54.46 55.22 +0.76
Acc Norm $\uparrow$ 3-bit 71.00 70.51 -0.49
Acc $\uparrow$ 3-bit 69.91 70.18 +0.27
Acc Norm $\uparrow$ 4-bit 73.34 69.97 -3.37
Acc $\uparrow$ 4-bit 71.98 71.11 -0.87
Acc Norm $\uparrow$ 8-bit 73.83 70.67 -3.16
Acc $\uparrow$ 8-bit 73.72 71.27 -2.45
Sciq Acc Norm $\uparrow$ 2-bit 45.80 45.30 -0.50
Acc $\uparrow$ 2-bit 43.90 43.80 -0.10
Acc Norm $\uparrow$ 3-bit 82.20 78.30 -3.90
Acc $\uparrow$ 3-bit 88.00 85.50 -2.50
Acc Norm $\uparrow$ 4-bit 85.90 80.40 -5.50
Acc $\uparrow$ 4-bit 90.50 86.60 -3.90
Acc Norm $\uparrow$ 8-bit 86.40 80.40 -6.00
Acc $\uparrow$ 8-bit 90.00 87.20 -2.80
Wikitext Word Perplexity $\downarrow$ 2-bit 4660.08 336.22 -4323.86
3-bit 19.24 21.77 +2.52
4-bit 15.84 18.46 +2.62
8-bit 15.10 17.86 +2.77
Table 13: Comparison of softpick vs softmax performance for HQQ [2] quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Quantization Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 2-bit 67.21 61.99 -5.22
Acc $\uparrow$ 2-bit 72.64 68.60 -4.04
Acc Norm $\uparrow$ 3-bit 67.21 61.99 -5.22
Acc $\uparrow$ 3-bit 72.64 68.60 -4.04
Lambada Acc $\uparrow$ 2-bit 49.56 43.61 -5.96
Perplexity $\downarrow$ 2-bit 11.38 15.83 +4.45
Acc $\uparrow$ 3-bit 49.56 43.61 -5.96
Perplexity $\downarrow$ 3-bit 11.38 15.83 +4.45
Piqa Acc Norm $\uparrow$ 2-bit 73.45 70.89 -2.56
Acc $\uparrow$ 2-bit 73.78 71.38 -2.39
Acc Norm $\uparrow$ 3-bit 73.45 70.89 -2.56
Acc $\uparrow$ 3-bit 73.78 71.38 -2.39
Sciq Acc Norm $\uparrow$ 2-bit 86.30 80.40 -5.90
Acc $\uparrow$ 2-bit 90.20 87.10 -3.10
Acc Norm $\uparrow$ 3-bit 86.30 80.40 -5.90
Acc $\uparrow$ 3-bit 90.20 87.10 -3.10
Wikitext Word Perplexity $\downarrow$ 2-bit 15.10 17.86 +2.77
3-bit 15.10 17.86 +2.77
Appendix E More Attention Maps
@@
Figure 8: More attention maps of the softmax 340M model on sample text input 1.
Figure 9: More attention maps of the softpick 340M model on sample text input 1.
Figure 10: More attention maps of the softmax 340M model on sample text input 2.
Figure 11: More attention maps of the softpick 340M model on sample text input 2.
Figure 12: More attention maps of the softmax 340M model on sample text input 3.
Figure 13: More attention maps of the softpick 340M model on sample text input 3.
Figure 14: More attention maps of the softmax 1.8B model on sample text input 1.
Figure 15: More attention maps of the softpick 1.8B model on sample text input 1.
Figure 16: More attention maps of the softmax 1.8B model on sample text input 2.
Figure 17: More attention maps of the softpick 1.8B model on sample text input 2.
Figure 18: More attention maps of the softmax 1.8B model on sample text input 3.
Figure 19: More attention maps of the softpick 1.8B model on sample text input 3.